OpenStudy (anonymous):

Rewrite in y=a(x-h)^2 by completing the square. y=x^2+8x-7. Thanks c:

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

I edited it.

OpenStudy (anonymous):

y = x² + 8x - 7 From completing the square formula : y = [x + (8/2) ]² - (8/2)² - 7 y = ( x + 4 )² - 16 - 7 y = ( x + 4 )² - 23

OpenStudy (anonymous):

Look at the following expansion (think in reverse):\[a(x-h)^2=a(x-2hx+h^2)=ax-2hx+ah^2\]Now take analogy to the equation you're given, and find \(a\) and \(h\).

OpenStudy (anonymous):

Correction: \(a(x-h)^2=a(x-2hx+h^2)=\mathbf{ax-2hax+ah^2}\)

OpenStudy (anonymous):

Sorry, real correction: \(a(x-h)^2=a(x-2hx+h^2)=\mathbf{ax^2-2hax+ah^2}\)

OpenStudy (lalaly):

completing the square\[y=x^2+8x-7\]\[y=(x+\frac{8}{2})^2-(\frac{8}{2})^2-7\]\[y=(x+4)^2-16-7\]\[y=(x+4)^2-23\]

OpenStudy (anonymous):

Thanks :)

OpenStudy (lalaly):

when u have \[y=x^2+bx+c\]completing the square u get\[y=(x+\frac{b}{2})^2-(\frac{b}{2})^2+c\]